OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL Saturn is 9.54 AU from the sun. What is its distance from the sun in Kilometers? A. 14,000,000 KM B. 15,723,270 km C. 57,000,000 km D. 1,431,000,000 km

OpenStudy (anonymous):

\[\color{green}{1~\text{AU} \approx 1.50 \times 10^8 ~\text{km} \\ 9.54~\text{AU} \approx 9.54 \times 1.50 \times 10^8~\text{km}\\\approx 1.43 \times 10^9~\text{km}}\]